Find the area of the region bounded y = 2x + 4, y = 1 and y = 3
and y-axis.

Grade:12

1 Answers

Rathod Shankar AskiitiansExpert-IITB
13 years ago

we have to find area bounded between lines y=1 and y=3 and y=2x+4 AND Y-AXIS

y=1 cuts the curve y=2x+4 at the point (-3/2, 1)

y=3 cuts the curve y=2x+4 at the point (-1/2, 3)

when we draw graph for these four curves thers ian traingle and rectangle. we have to compute those two surface areas.

 

reactangle has sides (0,1) (0,3) (-1/2,1), (-1/2,3)

so area = 2*1/2 =1

vortices of triangle (-3/2,1)  (-1/2,1), (-1/2,3)

area of traiangle= 2*1*1/2 =1

total area =2 units